OpenStudy (gabylovesyou):

Y- intercept is 16 ? x j(x) -2 -2 2 10 6 22

OpenStudy (fibonaccichick666):

seems unlikely, how did you get that?

OpenStudy (fibonaccichick666):

ok, let me double check your slope and answer. Give me a moment

OpenStudy (gabylovesyou):

ok ty :)

OpenStudy (phi):

that was close y-10= 3(x-2) y-10= 3x-6

OpenStudy (phi):

finish by adding 10 to both sides

OpenStudy (phi):

-6+10 is not 16

OpenStudy (fibonaccichick666):

\[m=\frac{10-(-2)}{2-(-2)}=\frac{12}{4}=3\] check

OpenStudy (fibonaccichick666):

and looks like phi got the other half

OpenStudy (gabylovesyou):

ohhhh i added 6 + 10

OpenStudy (phi):

watching signs is a pain, but necessary

OpenStudy (gabylovesyou):

ty :)

OpenStudy (gabylovesyou):

can u guys delete ur response... i did it in my own words but just in case -.-